40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 20 23:31:44 2011, Suresh, Update, IOO, MC Trans QPD is down 
    Reply  Thu Jul 21 12:19:29 2011, Jamie, Kiwamu, Update, IOO, MC Trans QPD working, now locking 
Message ID: 5009     Entry time: Wed Jul 20 23:31:44 2011     Reply to this: 5012
Author: Suresh 
Type: Update 
Category: IOO 
Subject: MC Trans QPD is down 

 

The mode cleaner is not locking because the MC Trans QPD signal is not present.  There is light on the QPD when the MC flashes and its position has not shifted.  The cable is plugged in well into the sensor head.  The signal cable is labled "MC2 Opt Lever"  and it arrives on the 1X4 rack along with other Optical Lever cables. Pressing the connector in did not solve the problem.

 

ELOG V3.1.3-